# TokenSelect: Efficient Long-Context Inference and Length Extrapolation for LLMs via Dynamic Token-Level KV Cache Selection

### v1

@@ 6.1. Experimental Settings
Implementation details.
[…]
6.2. Performance Comparisons
InfiniteBench.
As shown in Table 1, our TokenSelect achieves significantly superior overall performance on InfiniteBench compared to all baseline methods, even though TokenSelect uses the smallest token budget (¡3K). The fact that it significantly outperforms the original models demonstrates TokenSelect’s strong length extrapolation capability. We analyze that this is due to our adoption of a fine-grained KV Cache selection strategy, while considering the equal contribution of each head to selection, which ensures that we can select most critical tokens. Observing the performance of other methods, we find that RoPE interpolation methods (NTK, SelfExtend) generally perform poorly unless used on specially trained models such as Qwen2-7B-Instruct. The better performance of Qwen2-7B-Instruct on the original model can also be attributed to this. The sparse attention method StreamingLLM, based on fixed sparse patterns, can guarantee some of the model’s capabilities, but due to discarding a large amount of long-context information, it performs poorly on retrieval-related tasks (R.PK, R.Num, R.KV). The block-level selection method InfLLM can retain more long-context information compared to StreamingLLM. However, due to its sub-optimal block-level selection, it results in lower performance on most tasks compared to TokenSelect, even though we set a larger token budget for InfLLM. It is worth noting that Yi-1.5-6B does not perform normally on the R.KV task, as it is unable to correctly recite strings like the Universally Unique Identifier.
@@ 6.2. Performance Comparisons
Figure 6. Performance and Cache Rate with different similarity threshold $\theta$ of the Selection Cache on Qwen2-7B-Instruct.
RULER
To further demonstrate the long-context capability of TokenSelect, we conduct evaluation on the more challenging long-context benchmark RULER. Considering the increased difficulty of RULER and its substantial computational requirements, we include only comparable baseline methods. As shown in Table 2, our TokenSelect maintains significantly superior overall performance compared to other long-context inference methods. For all models, TokenSelect achieves length extrapolation while preserving the model’s original capabilities, benefiting from our efficient utilization of the model’s limited context length. Notably, due to the constraints of model’s context length, TokenSelect experiences performance degradation with larger token budgets (4K+4K) on Llama and Yi. However, its performance with smaller token budgets still significantly surpasses other baseline methods.
LongBench.
Due to space constraints, the results of LongBench are presented in the Appendix C. Although its relatively shorter text length makes it less suitable for evaluating state-of-the-art long-context inference methods, our TokenSelect still demonstrates superior overall performance compared to most baseline methods.
Table 3. Comparison of different methods on post-trained models on InfiniteBench and LongBench.
Methods InfiniteBench LongBench
En.Sum En.QA Code.D Math.F R.KV Avg.
LLaMA-3-8B-262K 20.2 12.4 22.1 26.6 14.4 33.9
    + MInference 20.5 12.9 22.3 33.1 12.8 38.4
Ours (w/ LLaMA-8K) 26.9 19.3 27.4 28.2 40.0 44.0
Comparing to methods based-on post-trained model.
In Table 3, we present the performance of the post-trained model and long-context inference method (Jiang et al., 2024) based on it. It shows that even compared to length extrapolation methods requiring additional training, the training-free TokenSelect still exhibits superior performance on most tasks. Although Minference can improve the performance of the original model, it fails to reverse the negative impact of long-text post-training on shorter text tasks (LongBench).
Table 4. Ablation study of the Selection Function $\mathcal{S}$ on InfiniteBench using Qwen2-7B-Instruct.
$\mathcal{S}$ En.QA En.MC Code.D Math.F R.Num R.KV
$\mathcal{I}_{\text{topk}}$ 15.15 45.85 28.43 31.14 98.47 16.60
$\mathcal{I}_{\text{head-vote}}$ 17.01 45.85 28.68 30.86 100.00 22.40
$\mathcal{I}_{\text{head-soft-vote}}$ 18.86 48.47 30.20 32.57 100.00 86.60
6.3. Ablation Studies
Table 5. Performance vs. Number of selected tokens $k$ on InfiniteBench using Qwen2-7B-Instruct.
$k$ En.Sum En.QA En.Mc Math.F R.Num R.KV
128 21.23 10.46 41.48 18.00 100.00 13.40
256 22.01 11.66 41.92 19.71 100.00 20.00
512 21.60 13.31 40.17 21.71 100.00 45.60
1K 21.35 15.13 44.10 24.57 100.00 73.00
2K 22.62 18.86 48.47 32.57 100.00 86.60
4K 24.09 21.11 51.53 21.71 100.00 88.00
8K 25.32 22.93 58.52 23.71 100.00 85.40
16K 26.54 23.04 62.88 28.16 100.00 72.00
@@ 6.3. Ablation Studies
Figure 7. Computation time for single chunk prefill step on different KV Cache lengths using Qwen2-7B-Instruct. The vertical axis represents the number of attended tokens, where SDPA denotes full attention by Flashinfer (chunk size: 512).
Figure 8. End to end latency per sample with different methods on InfiniteBench using Qwen2-7B-Instruct.
6.4. Hyper-parameter Analysis
@@
As shown in Table 5, we fixed $n_{\text{local}}$ to a relatively small value ($512$) to compare the performance when selecting different numbers of tokens. First, we observe that even selecting a very small number of tokens (e.g., 128, 256), our TokenSelect still demonstrates very comparable performance. Then, as $k$ increases, the effectiveness of TokenSelect further improves, indicating that more moderately critical tokens also contribute to the retention of long-context information. Finally, we find that when $k$ is set to larger values (e.g., 16K), our TokenSelect shows significant improvements in most tasks, further advancing the performance landscape of long-context inference methods.
@@ 6.5. Efficiency Comparisons
Fig. 8 compares the end-to-end latency of TokenSelect, InfLLM, and standard attention across various tasks. TokenSelect significantly accelerates long-context inference in real-world scenarios, achieving a maximum speedup of 4.70×4.70$\times$4.70 × over standard attention and 2.28×2.28$\times$2.28 × over the SOTA long-context inference method. Moreover, TokenSelect demonstrates superior performance compared to both of them.
6.6. Scaling Beyond 1 Million Context Length
To further explore TokenSelect’s performance in extreme long-context scenarios, we design an extended benchmark with different text lengths following InfiniteBench. As illustrated in the Fig. 9, our TokenSelect demonstrates the ability to accurately capture critical information with a small token budget in contexts up to 2M tokens, underscoring its potential in more application scenarios.
Figure 9. Performance comparison on extended R.PK and R.KV with different text length using Qwen2-7B-Instruct.
7. Conclusion
In this paper, we introduces TokenSelect, a model-agnostic and training-free approach for efficient long-context inference and length extrapolation. TokenSelect addresses the two major challenges faced by LLMs in processing long texts: the context length limitation from pre-training and the computational complexity of attention. This is achieved through a novel token-level selective sparse attention mechanism. Experimental results demonstrate that TokenSelect can achieve up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end inference latency, while exhibiting superior performance across multiple long-context benchmarks. This approach significantly enhances LLMs’ capability to handle long contexts, paving the way for efficient long-text processing in advancing Web applications.
Appendix A Formal Statement and Proof of Lemma 1
@@
$=(\hat{\mathbf{q}}_{1}\mathbf{k}_{p}\cos{\theta}-\hat{\mathbf{q}}_{1}\mathbf{k}_{p+1}\cos{\theta})-2\|\mathbf{k}\|_{\text{max}}\sin{\theta}$
(11) $\geq\hat{\eta}\cos{\theta}-2\|\mathbf{k}\|_{\text{max}}\sin{\theta}.$
@@
Appendix B More Information on Dataset and Metrics
For InfiniteBench (Zhang et al., 2024), we use longbook_sum_eng (En.Sum), longbook_qa_eng (En.QA), longbook_choice_eng (En.MC), longdialogue_qa_eng (En.Dia), code_debug (Code.D), math_find (Math.F), passkey (R.PK), number_string (R.Num) and kv_retrieval (R.KV) as evaluation datasets. The corresponding evaluation metrics are shown in Table 6. RULER (Hsieh et al., 2024) consists of various evaluation tasks: Single NIAH (needle in a haystack), Multi-keys NIAH, Multi-values NIAH, Multi-values NIAH, Multi-queries NIAH, Variable Tracking, Common Words Extraction, Frequent Words Extraction and Question Answering. The evaluation metric is match rate. For LongBench, we use all English tasks with evaluation metrics in Table 7.
Table 6. Evaluation metrics of different datasets on InfiniteBench.
@@ Appendix B More Information on Dataset and Metrics
Table 7. Evaluation metrics of different datasets on LongBench.
@@
Table 8. Comparison of different methods with different origin models on LongBench.
@@
Appendix C Experimental Results on LongBench
Compared to InfiniteBench and RULER, LongBench has much shorter text lengths. The 95% percentile for its lengths is 31K tokens. Considering that recent LLMs after SFT generally have context lengths of up to 32K tokens (Yang et al., 2024), LongBench is less suitable for evaluating state-of-the-art long-context inference methods. Nevertheless, as shown in Table 8, our TokenSelect still demonstrates superior overall performance compared to most baseline methods. It’s worth noting that Yi-1.5-6B did not yield effective results on the SAMSum task because it failed to correctly follow instructions.

### v2

@@ 6.2 Performance Comparisons
Table 2: Performance comparison on RULER.
RULER.
To further demonstrate the capability of TokenSelect, we conduct evaluation on the more challenging long-context benchmark RULER. Considering the increased difficulty of RULER and its substantial computational requirements, we include only comparable baseline methods. As shown in Table 2, our TokenSelect maintains significantly superior overall performance compared to other long-context inference methods. For all models, TokenSelect achieves length extrapolation while preserving the model’s original capabilities, benefiting from our efficient utilization of the model’s limited context length. Notably, due to the constraints of model’s context length, TokenSelect experiences performance degradation with larger token budgets (4K+4K) on Llama and Yi. However, its performance with smaller token budgets still significantly surpasses other baseline methods.
LongBench.
Due to space constraints, the results of LongBench are presented in the Appendix E. Although its relatively shorter text length makes it less suitable for evaluating state-of-the-art long-context inference methods, our TokenSelect still demonstrates superior overall performance compared to most baseline methods.
Methods InfiniteBench LongBench
En.Sum En.QA Code.D Math.F R.KV Avg.
Llama-3-8B-262K 20.2 12.4 22.1 26.6 14.4 33.9
    + MInference 20.5 12.9 22.3 33.1 12.8 38.4
Ours (w/ Llama-8K) 26.9 21.3 27.4 28.2 48.4 44.0
Table 3: Comparison of different methods on post-trained models on InfiniteBench and LongBench.
Comparing to methods based on post-trained model.
In Table 3, we present the performance of the post-trained model and long-context inference method Jiang et al. (2024) based on it. It shows that even compared to length extrapolation methods requiring additional training, the training-free TokenSelect still exhibits superior performance on most tasks. Although Minference can improve the performance of the original model, it fails to reverse the negative impact of long-text post-training on shorter text tasks (LongBench).
$\mathcal{S}$ En.QA En.MC Code.D R.PK R.Num R.KV
$\mathcal{I}_{\text{topk}}$ 15.15 45.85 28.43 100.00 98.47 16.60
$\mathcal{I}_{\text{head-vote}}$ 17.01 45.85 28.68 100.00 100.00 22.40
$\mathcal{I}_{\text{head-soft-vote}}$ 18.86 54.31 30.20 100.00 100.00 86.60
Table 4: Ablation study of the Selection Function $\mathcal{S}$ on InfiniteBench using Qwen2-7B-Instruct.
6.3 Ablation Studies
$k$ En.Sum En.QA En.Mc Math.F R.Num R.KV
128 21.23 10.46 41.48 18.00 100.00 13.40
256 22.01 11.66 41.92 19.71 100.00 20.00
512 21.60 13.31 40.17 21.71 100.00 45.60
1K 21.35 15.13 44.10 21.71 100.00 73.00
2K 22.62 18.86 54.31 21.71 100.00 86.60
4K 24.09 21.11 51.53 21.71 100.00 88.00
8K 25.32 22.93 58.52 23.71 100.00 85.40
16K 26.54 23.04 62.88 28.16 100.00 72.00
Table 5: Performance vs. Number of selected tokens $k$ on InfiniteBench using Qwen2-7B-Instruct.
@@ 6.3 Ablation Studies
Figure 6: Performance and Cache Rate with different threshold $\theta$ of the Selection Cache on Qwen2-7B-Instruct.
@@
Figure 7: Computation time vs. KV Cache lengths for single chunk prefill step using Qwen2-7B-Instruct. The vertical axis represents the number of attended tokens. SDPA denotes full attention by Flashinfer (chunk size: 512).
Figure 8: End to end latency per sample with different methods on InfiniteBench using Qwen2-7B-Instruct.
6.4 Hyper-parameter Analysis
@@
As shown in Table 5, we fix $n_{\text{local}}$ to a small value ($512$) to compare the performance when selecting different numbers of tokens. First, we observe that even selecting a very small number of tokens (e.g., 128, 256), our TokenSelect still demonstrates very comparable performance. Then, as $k$ increases, the effectiveness of TokenSelect further improves, indicating that more moderately critical tokens also contribute to the retention of long-context information. Finally, we find that when $k$ is set to larger values (e.g., 16K), our TokenSelect shows significant improvements in most tasks, further advancing the performance landscape of long-context inference methods.
@@ 6.5 Efficiency Comparisons
Fig. 8 compares the end-to-end latency of TokenSelect, InfLLM, and SDPA across various tasks. TokenSelect significantly accelerates long-context inference in real-world scenarios, achieving a maximum speedup of 4.70×4.70$\times$4.70 × over SDPA and 2.28×2.28$\times$2.28 × over the state-of-the-art long-context inference method while also delivering superior overall performance.
6.6 Scaling Beyond 1 Million Context Length
To further explore TokenSelect’s performance in extreme long-context scenarios, we design an extended benchmark with different text lengths following InfiniteBench. As illustrated in Fig. 9, our TokenSelect demonstrates the ability to accurately capture critical information with a small token budget in contexts up to 2M tokens, underscoring its potential in more application scenarios.
Figure 9: Performance comparison on extended R.PK and R.KV using Qwen2-7B-Instruct.
7 Conclusion
In this paper, we introduce TokenSelect, a training-free approach for efficient long-context inference and length extrapolation. TokenSelect addresses the two major challenges faced by LLMs in processing long texts: the context length limitation from pre-training and the computational complexity of attention. This is achieved through a novel token-level selective sparse attention mechanism. Experimental results demonstrate that TokenSelect can achieve up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end inference latency, while exhibiting superior performance across multiple long-context benchmarks.
8 Limitations
Our approach has inherent limitations that present opportunities for future work. A primary limitation of our method is that its training-free design—a significant advantage—acts as a double-edged sword, as its absolute performance is inherently tied to the quality of the underlying LLMs. Although our experiments demonstrate robustness of TokenSelect across various LLMs, some inherent shortcomings—such as the misrecognition of UUID strings by Yi-1.5-6B-Chat—indicate that certain issues may still require training to resolve. Moreover, while our method currently achieves state-of-the-art performance in long-context inference, recent long-text post-training techniques in the LLM community have shown impressive performance; notably, our TokenSelect is orthogonal to these approaches and can be employed during inference to trade a slight performance drop for significant efficiency gains. Finally, although our method achieves state-of-the-art efficiency improvements in long-context inference, the task remains inherently resource-intensive. For instance, even with an 8B-parameter model, complex benchmarks (e.g., RULER) can require approximately 8$\times$×A100 GPUs for nearly one day of runtime, and the computational cost is expected to increase substantially for larger models. We hope that our work, together with the community’s advances in model design, algorithm development, and infrastructure optimization, will help pave the way for further mitigating these computational challenges.
Appendix A Formal Statement and Proof of Lemma
@@
$=(\hat{\mathbf{q}}_{1}\mathbf{k}_{p}\cos{\theta}-\hat{\mathbf{q}}_{1}\mathbf{k}_{p+1}\cos{\theta})$
$\quad\quad\quad\quad\quad\quad-2\|\mathbf{k}\|_{\text{max}}\sin{\theta}$
$\geq\hat{\eta}\cos{\theta}-2\|\mathbf{k}\|_{\text{max}}\sin{\theta}.$ (11)
@@
Appendix B Detailed Descriptions on Baselines
In this paper, we use the following baselines:
  • NTK-Aware Scaled RoPE bloc97 (2023): A nonlinear RoPE interpolation method.
  • SelfExtend: A RoPE interpolation method that reuses the position ids of neighboring tokens.
  • StreamingLLM Xiao et al. (2024b): The state-of-the-ar method for long-context inference with predefined sparse patterns. Similar approaches include LM-Infinite Han et al. (2024).
  • InfLLM Xiao et al. (2024a): The state-of-the-art method for long-context inference and length extrapolation using a block-level selective sparse attention method.
  • MInference Jiang et al. (2024): The state-of-the-ar method for long-context prefilling acceleration, utilizing three sparse patterns including block-level sparse attention.
Appendix C More Information on Datasets
In this paper, we use the following datasets:
  • InfiniteBench Zhang et al. (2024a): The mainstream long-context benchmark consisting of multi-tasks. The average length of it exceeds 200K tokens.
  • RULER Hsieh et al. (2024): A challenging long-context benchmark containing 13 different tasks, with subsets of varying lengths up to 128K tokens.
  • LongBench Bai et al. (2024): Another mainstream long-context benchmark comprising 6 types of tasks. The 95% percentile for its lengths is 31K tokens.
For InfiniteBench Zhang et al. (2024a), we use longbook_sum_eng (En.Sum), longbook_qa_eng (En.QA), longbook_choice_eng (En.MC), longdialogue_qa_eng (En.Dia), code_debug (Code.D), math_find (Math.F), passkey (R.PK), number_string (R.Num) and kv_retrieval (R.KV) as evaluation datasets. The corresponding evaluation metrics are shown in Table 7. RULER Hsieh et al. (2024) consists of various evaluation tasks: Single NIAH (needle in a haystack), Multi-keys NIAH, Multi-values NIAH, Multi-values NIAH, Multi-queries NIAH, Variable Tracking, Common Words Extraction, Frequent Words Extraction and Question Answering. The evaluation metric is match rate. For LongBench, we use all English tasks with evaluation metrics in Table 8.
Appendix D Effiency Comparison with MInference
We note that Minference Jiang et al. (2024) has gained widespread adoption in real-world long-context inference applications due to its novel design of attention sparse patterns and efficient implementation based on vLLM. In the main text, we demonstrated TokenSelect’s performance advantages. To further prove its efficiency readiness for real-world applications, we followed Minference’s approach by comparing the end-to-end prefill latency under paged KV Cache management for different input token lengths on Llama-3-8B using a single A100, with results shown in Table 6. The results indicate that TokenSelect demonstrates significant advantages with shorter input token lengths, while maintaining efficiency comparable to MInference as input token lengths increase.
Length FlashAttention-2 (vLLM) MInference (vLLM) TokenSelect
1K 0.081 3.017 0.092
10K 0.832 2.762 1.290
50K 7.717 7.540 5.712
100K 21.731 14.081 12.088
128K 32.863 18.827 15.920
200K OOM OOM 26.500
300K OOM OOM 43.406
Table 6: Comparison of end-to-end prefill latency (s).
Appendix E Experimental Results on LongBench
Compared to InfiniteBench and RULER, LongBench has much shorter text lengths. The 95% percentile for its lengths is 31K tokens. Considering that recent LLMs after SFT generally have context lengths of up to 32K tokens Yang et al. (2024a), LongBench is less suitable for evaluating state-of-the-art long-context inference methods. Nevertheless, as shown in Table 9, our TokenSelect still demonstrates superior overall performance compared to most baseline methods. It’s worth noting that Yi-1.5-6B did not yield effective results on the SAMSum task because it failed to correctly follow instructions.
@@ Appendix E Experimental Results on LongBench
Table 7: Evaluation metrics of different datasets on InfiniteBench.
@@
Table 8: Evaluation metrics of different datasets on LongBench.
@@
Table 9: Comparison of different methods with different origin models on LongBench.